Train A has a speed 15mph faster than train B. I the same time that train A can travel150 miles, train B travels 105miles. Find the rate of both trains.
Answer by Alan3354(69443)

Train A has a speed 15mph faster than train B. I the same time that train A can travel 150 miles, train B travels 105miles. Find the rate of both trains.
---------------
A = 150/t
B = 105/t
A = B + 15
150/t = 105/t + 15
150 = 105 + 15t
t = 3 hours
-------------
A = 50
B = 35